# Confidential Databases Without Cryptographic Mappings

### 19 Mar 2026 v1

@@ 1 Introduction
Cloud databases have become the backbone of modern digital infrastructure, with organizations increasingly outsourcing their data management to Database-as-a-Service (DBaaS) providers amazon-aurora; azure-sql; google-sql. However, this convenience comes with a critical security risk: cloud administrators have unrestricted access to all tenant data, leaving opportunities for both malicious and accidental exposure. Recent breaches, such as the 2025 Allianz Life exposure of 1.1 million personal records data-breach-2025, underscore this threat. For sensitive databases regulated by laws such as HIPAA hippa and PCI-DSS pci-dss, entrusting cloud infrastructure with plaintext data is untenable.
Confidential databases (CDBs) have emerged as a compelling solution. By leveraging trusted execution environments (TEEs), CDBs enable secure data upload and query execution while ensuring data remains encrypted from the cloud provider. Over the past two decades, researchers have explored various CDB designs Bajaj2011TrustedDB; Arasu2013Cipherbase; Priebe2018EnclaveDB; Vinayagamurthy2019StealthDB; Fuhry2020EncDBDB, and major cloud providers have deployed CDB systems Antonopoulos2020AEv2; Wang2022Operon; Guo2021GaussDB; Li2023HEDB; yang2024secudb.
Despite this industry momentum, a critical roadblock to widespread adoption of CDB systems remains: severe performance degradation compared to plaintext databases, with analytical queries experiencing up to 79.5 higher latency Li2023HEDB. This gap arises from a deliberate system design choice made by industry deployments: rather than enclosing the entire DBMS within a TEE, which would prevent database administrators from performing essential maintenance tasks, modern CDBs execute only critical operators (e.g., numeric SUM operations) inside TEEs. The resulting split architecture necessitates synchronous software-based decryption and encryption for every remote procedure call (RPC) across domains, as illustrated in Figure 1. The computational burden scales linearly with data volume. For instance, a SUM aggregation over an -row table requires 2-2 decryptions and -1 encryptions in total. Even worse, in order to support arbitrary computation across fields, modern CDBs employ field-level encryption, inflating each 4-byte integer to 32 bytes (8 expansion) due to cryptographic metadata (12-byte nonce and 16-byte authentication tag required by schemes like AES-GCM). This dramatically amplifies cross-domain data movement, memory pressure, and I/O bandwidth demands. These bottlenecks raise a fundamental question: Is cryptographic overhead inherent to CDB design, or can we achieve the same security guarantees with far cheaper alternatives?
@@
We evaluate Fedb against HEDB Li2023HEDB, a state-of-the-art modern CDB that achieves the trifecta of functionality, maintainability, and security. Our results show that Fedb achieves substantial performance improvements while preserving all desired properties. On the transactional benchmark TPC-C, Fedb achieves up to 1.8 speedup over HEDB. On the analytical benchmark TPC-H, Fedb reduces query execution time by up to 78.0. Fedb outperforms HEDB with 2.9-4.9 speedups. Additionally, Fedb reduces storage overhead by 80.0% and 34.3% across two workloads, respectively. Fedb has been open-sourced and integrated into production GaussDB systems111https://gitee.com/openeuler/secGear/pulls/354.
@@ 2.1 Confidential Computing
Confidential computing amd-sev; intel-sgx; intel-tdx; arm-cca; aws-nitro; lee2020keystone; feng2021penglai; Li2021TwinVisor; Hunt2021PEF; Li2022CCA provides hardware-based trusted execution environments (TEEs) with three security guarantees. First, secure isolation through either enclaves intel-sgx; feng2021penglai; lee2020keystone or confidential virtual machines amd-sev; intel-tdx; arm-cca protects against inspection or tampering by untrusted privileged software. Second, hardware root-of-trust enables clients to verify code authenticity via remote attestation. Third, memory encryption and optional integrity checks resist physical attacks. Confidential computing is now widely available on commercial clouds google-sev; azure-sgx; aws-nitro; ibm-hyperprotect; alibaba-cc.
@@ 2.2 Confidential Databases
Database-as-a-Service (DBaaS) Hacigumus2002DBaaS offers fully managed databases that handle updates, backups, and maintenance for users. However, since cloud environments are inherently untrusted, sensitive data like healthcare and financial information requires secure storage and processing.
Confidential databases (CDBs) Bajaj2011TrustedDB; Arasu2013Cipherbase; Priebe2018EnclaveDB; Ribeiro2018DBStore; Vinayagamurthy2019StealthDB; Fuhry2020EncDBDB; Li2023HEDB; Antonopoulos2020AEv2; Guo2021GaussDB; Wang2022Operon; yang2024secudb address this by leveraging confidential computing to enable SQL query processing over encrypted data without compromising confidentiality. Industry-deployed CDBs Antonopoulos2020AEv2; Guo2021GaussDB; Wang2022Operon; Li2023HEDB adopt a split architecture: only essential expression operators (e.g., arithmetic, comparison222All operators yield encrypted results, except for comparisons (e.g. , , ), which output plaintext boolean values for use in indexing ciphertexts.) execute inside TEEs, while most DBMS logic (e.g., SQL parsing, planning, buffer management) remains in the untrusted domain. This minimizes the TCB and preserves maintainability: DBAs can inspect query plans and debug DBMS issues. We refer to these industry-adopted systems as “modern CDBs.”
@@
HEDB Li2023HEDB: a representative modern CDB. Modern CDBs could be vulnerable to smuggle attacks Li2023HEDB: malicious DBAs exploit operator interfaces to construct known ciphers (via arithmetic operators, e.g., division yields encrypted “one,” addition builds sequences) and compare them against victim ciphertexts (via equality operators), thereby extracting secrets without authorization. HEDB Li2023HEDB addresses smuggle attacks through a dual-TEE architecture: an integrity zone that hosts the DBMS and a privacy zone that hosts operators, with a secure channel preventing direct interface access by DBAs. For maintainability, HEDB records cross-zone calls in ciphertext, enabling DBAs to replay and debug queries without plaintext access. Fedb inherits and improves this design, but is designed to work with any modern CDB.
@@ 2.3 What Makes Modern CDB Systems Slow?
Modern CDBs exemplified by HEDB Li2023HEDB have demonstrated significant performance overheads. Benchmarks on TPC-H show that HEDB can incur up to a 79.5 increase in latency compared to a plaintext database. These overheads render modern CDBs impractical for latency-sensitive applications, hindering broader industry adoption. Our analysis attributes them to two main factors.
@@
Empirical analysis. We profile HEDB Li2023HEDB (commit f48742de) using three workloads: (i) TPC-H Q1 with a dataset larger than memory, representing large table computation; (ii) TPC-H Q1 with a dataset fitting in memory, capturing lightweight analytical settings; and (iii) a transactional benchmark with many write-heavy small queries. Figure 2 presents the breakdown of execution latency and storage costs. Despite the batching optimization, HEDB exhibits significantly higher overhead compared to the plaintext database. As shown in Figure 2 (a), the en/decryption operations, along with I/O amplification overhead, contribute considerably to the overall execution time, accounting for 10.5-62.6% and 13.9-25.5% across different workloads. Combined, they cause an overall latency slowdown of up to 5.3 compared to plaintext. Figure 2 (b) also shows that the field-level encryption incurs 1.4 to 3 storage overhead compared to plaintext. The remaining 8.6-26.5% of the execution time is attributable to DBMS execution and RPC overhead across domains, the latter of which stems from frequent copying of expanded ciphertexts.
Strawman: a ciphertext-to-plaintext cache. Since confidential VM-based TEEs (e.g., Intel TDX intel-tdx, AMD SEV-SNP amd-sev) provide sufficiently large secure memory, one can maintain a ciphertext-to-plaintext cache within TEEs to eliminate cryptographic operations from the critical path. However, this has two drawbacks. First, ciphertext search costs can be substantial: our preliminary experiment using C++’s STL unordered_map with one million ciphertexts showed an average lookup time of 1,601 CPU cycles, exceeding the constant-time decryption cost of 1,408 cycles. Second, storing both ciphertexts and plaintexts does not address the overhead of ciphertext expansion.
@@ 3 Fedb Overview
Threat model. We assume a powerful adversary who controls the entire cloud infrastructure except TEE hardware. This includes compromised server software, rogue cloud and database administrators, and external attackers. Adversaries can read and modify any data outside TEE domains—data at rest, in untrusted memory, and in transit—and observe all network traffic and I/O patterns between TEE memory and persistent storage. We do not consider database vulnerabilities dbms-bugs, TEE defects Kocher2019Spectre, denial-of-service or physical attacks.
@@
In fact, separating indirection from protection is a long-standing systems principle lampson2020hints. Capability systems dennis1980capability; levy2014capability used unforgeable tokens protected by OS confinement, not cryptography. Similarly, OSes have long used opaque handles, e.g., Unix file descriptors ritchie1974unix; accetta1986mach, that are mapped internally by the kernel. In both cases, security comes from address-space isolation, keeping references simple and fast.
@@
With minor modifications using the provided APIs, existing modern CDBs Vinayagamurthy2019StealthDB; Antonopoulos2020AEv2; Guo2021GaussDB; Wang2022Operon; Li2023HEDB can enjoy the benefits of Fedb’s performance gains (see 1).
Workflow. Fedb’s query execution follows modern CDBs (described in § 2.2). To perform computation, the proxy fetches secrets associated with each input FID from the mapping store via Get(), invokes the appropriate expression operator to compute over these retrieved secrets, and then stores the result back to the mapping store via Put(), which generates a new FID representing the output. The DBMS proceeds with query execution using this returned FID. Finally, the proxy translates FIDs back into ciphertexts by calling Get() followed by encrypting the retrieved results within the privacy zone before returning them to users. This design preserves compatibility (G1) with modern CDBs that employ client-side encryption Antonopoulos2020AEv2; Guo2021GaussDB; Wang2022Operon; Li2023HEDB.
@@ 4.1 Mapping Management
Data structure design. The mapping store is a linear array of tightly packed fields. Each sensitive field is assigned a unique integer as the FID that serves as a direct array index, hiding the size of individual fields from the adversary. This design restricts the metadata overhead per sensitive field to a mere 8B (FID), representing a 75% reduction compared to the 32B metadata required by modern CDBs. Compared to tree- or hash-based key-value stores (e.g., LevelDB leveldb, RocksDB rocksdb, LMDB lmdb), the design provides two advantages. First, it enables access complexity to sensitive fields in the privacy zone, eliminating the need for tree traversals or hash table probing. Second, synchronization is only necessary during FID allocation, which significantly reduces lock contention.
Mapping store operations. For fixed-length types (e.g., FLOAT8, TIMESTAMP), the Delete() operation marks the corresponding mapping entries as logically deleted. When inserting with Put(), Fedb either overwrites logically deleted slots or appends the new data to the end. For variable-length types (e.g., VARCHAR, NUMERIC), the situation is subtle because values cannot be overwritten in place. We address this by maintaining two arrays within the privacy zone: an offset array and a slab-like data array. The offset array is indexed directly by FIDs and stores pointers to the actual field. The data array is inspired by the slab allocator bonwick1994slab: fields are grouped into size-class buckets, and each bucket is maintained as a contiguous packed array. This minimizes memory fragmentation for frequent allocation and deallocation. Operations within a single bucket follow the same overwrite/append logic as the fixed-length case.
@@ 5 Ensuring Cross-domain Consistency
We adopt external synchrony nightingale2008rethink, a consistency model that distinguishes between internal system state (invisible to users) and external state (user-visible). External synchrony requires only that the external state remains consistent—internal states may temporarily diverge as long as they never violate external invariants. In Fedb, the mapping store constitutes internal state (users never see it), while the FIDs stored in the database represent external state. Therefore, external synchrony requires maintaining the invariant that every FID in the database maps to a valid, up-to-date secret in the mapping store. Temporary "orphan" secrets in the mapping store without corresponding database FIDs are permissible, as they remain invisible to users. To uphold this invariant, Fedb introduces lightweight synchronization between the DBMS and mapping store at transaction boundaries. We now detail the mechanisms for handling aborts, commits, and crash recovery.
@@ 6 Implementation
We implemented Fedb in 5.8K lines of C++ code on PostgreSQL 15.5. Table 2 lists the code complexity of each component. Fedb leverages ARMv8.4 Secure EL2 armv8-manual; Li2021TwinVisor to create two isolated confidential VMs as integrity and privacy zones. Both use dm-integrity dm-integrity for corruption detection; the privacy zone additionally applies dm-crypt dm-crypt for transparent I/O encryption. The mapping store is implemented as memory-mapped (mmap-based) arrays for persistence. Temporal locality is handled by an LRU cache combined with Linux sequential prefetching, while spatial locality is achieved by aligning partitions with database tables. Each 64-bit FID encodes a configurable 16-bit partition prefix and a 48-bit offset; unique FIDs within a partition are generated via an atomic monotonic counter. Users can tune FID length and partition-prefix size according to expected data volume—shorter FIDs reduce storage overhead and improve lookup efficiency.
@@
| Operators (from HEDB Li2023HEDB) | C++ | |
@@ 7.1 End-to-end Benchmarks (RQ1)
We use two standard benchmarks: TPC-C and TPC-H. For comparison, we include the plaintext database (original PostgreSQL) and a state-of-the-art, HEDB Li2023HEDB.
TPC-C. Following Microsoft Always Encrypted Antonopoulos2020AEv2, we encrypt personally identifiable columns (e.g., C_STREET_1, C_CITY, C_STATE in the Customer table). We evaluate TPC-C throughput (transactions per second, or TPS) across varying numbers of connection threads using 128 warehouses. To prevent data-size leakage, we pad variable-length types (e.g., VARCHAR) to a fixed length of 128 B. After a 120-second warm-up, we measure for 600 seconds. As shown in Figure 5, throughput peaks at 8 connections and then degrades due to contention. Across all bars, Fedb consistently outperforms HEDB. HEDB achieves only 51.6–77.3% of baseline throughput, whereas Fedb reaches 63.3–91.3%, delivering up to 1.8 speedup.
TPC-H. Following Alibaba HEDB Li2023HEDB, we use TPC-H (scale factor = 3), a standard OLAP benchmark with large joins, group-bys, aggregations, and arithmetic operations. We make all non-key columns sensitive, as TPC-H does not define security attributes. Both Fedb and HEDB can enable batching (with a batch size of 256 fields), which aggregates multiple operations into a single RPC instead of iteratively. We also report results with batching disabled. Results are averaged over 5 runs per query after one warm-up run.

### 23 Mar 2026 v2

@@ 1 Introduction
Cloud databases have become the backbone of modern digital infrastructure, with organizations increasingly outsourcing their data management to Database-as-a-Service (DBaaS) providers [1, 2, 3]. However, this convenience comes with a critical security risk: cloud administrators have unrestricted access to all tenant data, leaving opportunities for both malicious and accidental exposure. Recent breaches, such as the 2025 Allianz Life exposure of 1.1 million personal records [4], underscore this threat. For sensitive databases regulated by laws such as HIPAA [5] and PCI-DSS [6], entrusting cloud infrastructure with plaintext data is untenable.
Confidential databases (CDBs) have emerged as a compelling solution. By leveraging trusted execution environments (TEEs), CDBs enable secure data upload and query execution while ensuring data remains encrypted from the cloud provider. Over the past two decades, researchers have explored various CDB designs [7, 8, 9, 10, 11], and major cloud providers have deployed CDB systems [12, 13, 14, 15, 16].
Despite this industry momentum, a critical roadblock to widespread adoption of CDB systems remains: severe performance degradation compared to plaintext databases, with analytical queries experiencing up to 79.5 higher latency [15]. This gap arises from a deliberate system design choice made by industry deployments: rather than enclosing the entire DBMS within a TEE, which would prevent database administrators from performing essential maintenance tasks, modern CDBs execute only critical operators (e.g., numeric SUM operations) inside TEEs. The resulting split architecture necessitates synchronous software-based decryption and encryption for every remote procedure call (RPC) across domains, as illustrated in Figure 1. The computational burden scales linearly with data volume. For instance, a SUM aggregation over an -row table requires 2-2 decryptions and -1 encryptions in total. Even worse, in order to support arbitrary computation across fields, modern CDBs employ field-level encryption, inflating each 4-byte integer to 32 bytes (8 expansion) due to cryptographic metadata (12-byte nonce and 16-byte authentication tag required by schemes like AES-GCM). This dramatically amplifies cross-domain data movement, memory pressure, and I/O bandwidth demands. These bottlenecks raise a fundamental question: Is cryptographic overhead inherent to CDB design, or can we achieve the same security guarantees with far cheaper alternatives?
@@
We evaluate Fedb against HEDB [15], a state-of-the-art modern CDB that achieves the trifecta of functionality, maintainability, and security. Our results show that Fedb achieves substantial performance improvements while preserving all desired properties. On the transactional benchmark TPC-C, Fedb achieves up to 1.8 speedup over HEDB. On the analytical benchmark TPC-H, Fedb reduces query execution time by up to 78.0. Fedb outperforms HEDB with 2.9-4.9 speedups. Additionally, Fedb reduces storage overhead by 80.0% and 34.3% across two workloads, respectively. Fedb has been open-sourced and integrated into production GaussDB systems111https://gitee.com/openeuler/secGear/pulls/354.
@@ 2.1 Confidential Computing
Confidential computing [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] provides hardware-based trusted execution environments (TEEs) with three security guarantees. First, secure isolation through either enclaves [18, 23, 22] or confidential virtual machines [17, 19, 20] protects against inspection or tampering by untrusted privileged software. Second, hardware root-of-trust enables clients to verify code authenticity via remote attestation. Third, memory encryption and optional integrity checks resist physical attacks. Confidential computing is now widely available on commercial clouds [27, 28, 21, 29, 30].
@@ 2.2 Confidential Databases
Database-as-a-Service (DBaaS) [31] offers fully managed databases that handle updates, backups, and maintenance for users. However, since cloud environments are inherently untrusted, sensitive data like healthcare and financial information requires secure storage and processing.
Confidential databases (CDBs) [7, 8, 9, 32, 10, 11, 15, 12, 14, 13, 16] address this by leveraging confidential computing to enable SQL query processing over encrypted data without compromising confidentiality. Industry-deployed CDBs [12, 14, 13, 15] adopt a split architecture: only essential expression operators (e.g., arithmetic, comparison222All operators yield encrypted results, except for comparisons (e.g. , , ), which output plaintext boolean values for use in indexing ciphertexts.) execute inside TEEs, while most DBMS logic (e.g., SQL parsing, planning, buffer management) remains in the untrusted domain. This minimizes the TCB and preserves maintainability: DBAs can inspect query plans and debug DBMS issues. We refer to these industry-adopted systems as “modern CDBs.”
@@
HEDB [15]: a representative modern CDB. Modern CDBs could be vulnerable to smuggle attacks [15]: malicious DBAs exploit operator interfaces to construct known ciphers (via arithmetic operators, e.g., division yields encrypted “one,” addition builds sequences) and compare them against victim ciphertexts (via equality operators), thereby extracting secrets without authorization. HEDB [15] addresses smuggle attacks through a dual-TEE architecture: an integrity zone that hosts the DBMS and a privacy zone that hosts operators, with a secure channel preventing direct interface access by DBAs. For maintainability, HEDB records cross-zone calls in ciphertext, enabling DBAs to replay and debug queries without plaintext access. Fedb inherits and improves this design, but is designed to work with any modern CDB.
@@ 2.3 What Makes Modern CDB Systems Slow?
Modern CDBs exemplified by HEDB [15] have demonstrated significant performance overheads. Benchmarks on TPC-H show that HEDB can incur up to a 79.5 increase in latency compared to a plaintext database. These overheads render modern CDBs impractical for latency-sensitive applications, hindering broader industry adoption. Our analysis attributes them to two main factors.
@@
Empirical analysis. We profile HEDB [15] (commit f48742de) using three workloads: (i) TPC-H Q1 with a dataset larger than memory, representing large table computation; (ii) TPC-H Q1 with a dataset fitting in memory, capturing lightweight analytical settings; and (iii) a transactional benchmark with many write-heavy small queries. Figure 2 presents the breakdown of execution latency and storage costs. Despite the batching optimization, HEDB exhibits significantly higher overhead compared to the plaintext database. As shown in Figure 2 (a), the en/decryption operations, along with I/O amplification overhead, contribute considerably to the overall execution time, accounting for 10.5-62.6% and 13.9-25.5% across different workloads. Combined, they cause an overall latency slowdown of up to 5.3 compared to plaintext. Figure 2 (b) also shows that the field-level encryption incurs 1.4 to 3 storage overhead compared to plaintext. The remaining 8.6-26.5% of the execution time is attributable to DBMS execution and RPC overhead across domains, the latter of which stems from frequent copying of expanded ciphertexts.
Strawman: a ciphertext-to-plaintext cache. Since confidential VM-based TEEs (e.g., Intel TDX [19], AMD SEV-SNP [17]) provide sufficiently large secure memory, one can maintain a ciphertext-to-plaintext cache within TEEs to eliminate cryptographic operations from the critical path. However, this has two drawbacks. First, ciphertext search costs can be substantial: our preliminary experiment using C++’s STL unordered_map with one million ciphertexts showed an average lookup time of 1,601 CPU cycles, exceeding the constant-time decryption cost of 1,408 cycles. Second, storing both ciphertexts and plaintexts does not address the overhead of ciphertext expansion.
@@ 3 Fedb Overview
Threat model. We assume a powerful adversary who controls the entire cloud infrastructure except TEE hardware. This includes compromised server software, rogue cloud and database administrators, and external attackers. Adversaries can read and modify any data outside TEE domains—data at rest, in untrusted memory, and in transit—and observe all network traffic and I/O patterns between TEE memory and persistent storage. We do not consider database vulnerabilities [33], TEE defects [34], denial-of-service or physical attacks.
@@
In fact, separating indirection from protection is a long-standing systems principle [35]. Capability systems [36, 37] used unforgeable tokens protected by OS confinement, not cryptography. Similarly, OSes have long used opaque handles, e.g., Unix file descriptors [38, 39], that are mapped internally by the kernel. In both cases, security comes from address-space isolation, keeping references simple and fast.
@@
With minor modifications using the provided APIs, existing modern CDBs [10, 12, 14, 13, 15] can enjoy the benefits of Fedb’s performance gains (see 1).
Workflow. Fedb’s query execution follows modern CDBs (described in § 2.2). To perform computation, the proxy fetches secrets associated with each input FID from the mapping store via Get(), invokes the appropriate expression operator to compute over these retrieved secrets, and then stores the result back to the mapping store via Put(), which generates a new FID representing the output. The DBMS proceeds with query execution using this returned FID. Finally, the proxy translates FIDs back into ciphertexts by calling Get() followed by encrypting the retrieved results within the privacy zone before returning them to users. This design preserves compatibility (G1) with modern CDBs that employ client-side encryption [12, 14, 13, 15].
@@ 4.1 Mapping Management
Data structure design. The mapping store is a linear array of tightly packed fields. Each sensitive field is assigned a unique integer as the FID that serves as a direct array index, hiding the size of individual fields from the adversary. This design restricts the metadata overhead per sensitive field to a mere 8B (FID), representing a 75% reduction compared to the 32B metadata required by modern CDBs. Compared to tree- or hash-based key-value stores (e.g., LevelDB [40], RocksDB [41], LMDB [42]), the design provides two advantages. First, it enables access complexity to sensitive fields in the privacy zone, eliminating the need for tree traversals or hash table probing. Second, synchronization is only necessary during FID allocation, which significantly reduces lock contention.
Mapping store operations. For fixed-length types (e.g., FLOAT8, TIMESTAMP), the Delete() operation marks the corresponding mapping entries as logically deleted. When inserting with Put(), Fedb either overwrites logically deleted slots or appends the new data to the end. For variable-length types (e.g., VARCHAR, NUMERIC), the situation is subtle because values cannot be overwritten in place. We address this by maintaining two arrays within the privacy zone: an offset array and a slab-like data array. The offset array is indexed directly by FIDs and stores pointers to the actual field. The data array is inspired by the slab allocator [43]: fields are grouped into size-class buckets, and each bucket is maintained as a contiguous packed array. This minimizes memory fragmentation for frequent allocation and deallocation. Operations within a single bucket follow the same overwrite/append logic as the fixed-length case.
@@ 5 Ensuring Cross-domain Consistency
We adopt external synchrony [44], a consistency model that distinguishes between internal system state (invisible to users) and external state (user-visible). External synchrony requires only that the external state remains consistent—internal states may temporarily diverge as long as they never violate external invariants. In Fedb, the mapping store constitutes internal state (users never see it), while the FIDs stored in the database represent external state. Therefore, external synchrony requires maintaining the invariant that every FID in the database maps to a valid, up-to-date secret in the mapping store. Temporary "orphan" secrets in the mapping store without corresponding database FIDs are permissible, as they remain invisible to users. To uphold this invariant, Fedb introduces lightweight synchronization between the DBMS and mapping store at transaction boundaries. We now detail the mechanisms for handling aborts, commits, and crash recovery.
@@ 6 Implementation
We implemented Fedb in 5.8K lines of C++ code on PostgreSQL 15.5. Table 2 lists the code complexity of each component. Fedb leverages ARMv8.4 Secure EL2 [45, 24] to create two isolated confidential VMs as integrity and privacy zones. Both use dm-integrity [46] for corruption detection; the privacy zone additionally applies dm-crypt [47] for transparent I/O encryption. The mapping store is implemented as memory-mapped (mmap-based) arrays for persistence. Temporal locality is handled by an LRU cache combined with Linux sequential prefetching, while spatial locality is achieved by aligning partitions with database tables. Each 64-bit FID encodes a configurable 16-bit partition prefix and a 48-bit offset; unique FIDs within a partition are generated via an atomic monotonic counter. Users can tune FID length and partition-prefix size according to expected data volume—shorter FIDs reduce storage overhead and improve lookup efficiency.
@@
| Operators (from HEDB [15]) | C++ | |
@@ 7.1 End-to-end Benchmarks (RQ1)
We use two standard benchmarks: TPC-C and TPC-H. For comparison, we include the plaintext database (original PostgreSQL) and a state-of-the-art, HEDB [15].
TPC-C. Following Microsoft Always Encrypted [12], we encrypt personally identifiable columns (e.g., C_STREET_1, C_CITY, C_STATE in the Customer table). We evaluate TPC-C throughput (transactions per second, or TPS) across varying numbers of connection threads using 128 warehouses. To prevent data-size leakage, we pad variable-length types (e.g., VARCHAR) to a fixed length of 128 B. After a 120-second warm-up, we measure for 600 seconds. As shown in Figure 5, throughput peaks at 8 connections and then degrades due to contention. Across all bars, Fedb consistently outperforms HEDB. HEDB achieves only 51.6–77.3% of baseline throughput, whereas Fedb reaches 63.3–91.3%, delivering up to 1.8 speedup.
TPC-H. Following Alibaba HEDB [15], we use TPC-H (scale factor = 3), a standard OLAP benchmark with large joins, group-bys, aggregations, and arithmetic operations. We make all non-key columns sensitive, as TPC-H does not define security attributes. Both Fedb and HEDB can enable batching (with a batch size of 256 fields), which aggregates multiple operations into a single RPC instead of iteratively. We also report results with batching disabled. Results are averaged over 5 runs per query after one warm-up run.